OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that passes through (-5,3) and is perpendicular to the line described by y = 5x.

OpenStudy (anonymous):

slope intercept form is y=mx+b now to get a line perpendicular to a line with a slope of 5 you need the reciprocal ..or -1/5 so y=-1/5x and now you need the intercept...its useful to have graph paper here... just start from the point given and use the slope you found...and you willl get an intercept of 2 ... so your equation is y=-1/5+2